Use f prime left parenthesis x right parenthesis equals ModifyingBelow lim With h right arrow 0 StartFraction f left parenthesis x plus h right parenthesis minus f left parenthesis x right parenthesis Over h EndFractionf′(x)=limh→0 f(x+h)−f(x) h to find the derivative at x for the given function. s left parenthesis x right parenthesis equals 2 x plus 6s(x)=2x+6 s prime left parenthesis x right parenthesiss′(x)equals=nothing
Answer #1

Solution:

s(x) = 2x+6

$ The derivative of function $ f(x) $ is given by:

f'(x) = \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h}

i.e., \ \ \ s'(x) = \lim_{h \rightarrow 0} \frac{s(x+h) - s(x)}{h}

i.e., \ \ \ s'(x) = \lim_{h \rightarrow 0} \frac{2(x+h) +6 - 2x -6}{h}

   = \lim_{h \rightarrow 0} \frac{2x+2h +6 - 2x -6}{h}

   = \lim_{h \rightarrow 0} \frac{2h }{h}

   = \lim_{h \rightarrow 0} (2)

   = 2

$ Therefore, the derivative of $ s(x) $ is: \ $ s'(x) = 2